OpenStudy (anonymous):

graph using the slope and the y intercept x+5y=20

OpenStudy (anonymous):

the eq is x+5y=20 we need to rearrange in the form y=mx+c 5y=-x+20 y=-x/5 + 4 m= 1/5 c= 4